OpenStudy (anonymous):

Find polar coordinates from the rectangular coordinates (10,3)

OpenStudy (anonymous):

Polar coordinate is in the form of (r,θ) where, 1)r^2=x^2+y^2 2)θ=arctan(y/x) Use these formulas to find r and θ. You have x=10 and y=3.

OpenStudy (anonymous):

so would be 9/100 so .09? and 109 as r^2 so squared 10.5?

OpenStudy (anonymous):

\[r=\sqrt{109}, \theta=\tan^{-1}({3 \over 10})\]

OpenStudy (anonymous):

Approximate answer is (10.4,0.29)

OpenStudy (anonymous):

omg thanks i have to get this done by tommorow or i wont get my navy nuke job mind helping me with few others i through up/

OpenStudy (anonymous):

np